OpenStudy (anonymous):

Explain why the following simulation fails to model the situation properly. Use a random integer from 0 to 3 to represent the number of girls in a family of 3 children. The answer is : The outcomes are not equally likely. The simulation assumes they are. Correct Reasoning: Having 0 through 3 girls in a family of 3 children are not equally likely. For example, having a total of 3 girls is less likely than having 2 girls out of 3 children. The simulation assumes they are equally likely. My question is why aren't the outcomes equally likely?

OpenStudy (whpalmer4):

for each child, it's approximately 50/50 whether it will be male or female. think of a coin labeled male on one side and female on the other. having one kid, you flip the coin once. having two kids, you flip the coin twice. here are the possible outcomes: male / male male / female female / male female / female 4 different outcomes: 1 outcome gives you 2 girls 1 outcome gives you 2 boys 2 outcomes give you a boy and a girl if we extend this to 3 kids: male/male / male male/male / female male/female / male male/female / female female/male / male female/male / female female/female / male female/female / female 8 outcomes: 1 outcome gives you 3 girls 1 outcome gives you 3 boys 3 outcomes give you 2 boys and a girl 3 outcomes give you 2 girls and a boy

OpenStudy (ybarrap):

Look at the sample space for 3 children. Let b = boy, g = girl. These are the possibilities (in order of age, youngest to oldest): bbb bbg bgb bgg gbb gbg ggb ggg Possibilities 0 girls = 1 1 girl = 3 2 girls = 2 3 girls = 1 Probabilities P(0 girls) = 1/8 P(1 girl) = 3/8 P(2 girls) = 3/8 P(3 girls) = 1/8 So now you see that the probability of 3 girls is less likely than the probability of 2 girls. In fact, having 2 girls is three times as likely than having 3 girls! Simply because there are more ways for this to happen. Therefore, the probability of 0 to 3 girls is NOT equally likely. You might wonder why age is important. Say we did not include it as a factor, then our sample space would look like this if age didn't matter: bbb bbg bgg ggg In this case you might think that having 0, 1, 2 or 3 girls would be equally likely. It's like counting the number of times you get 2 tails out of three. You can't just count the times this happens in the 2nd and 3rd flip, you also need to count the number of times it happens on the first. Otherwise, your sample space for coins would look like this: hhh hht htt ttt But you know in fact there are other ways to get 2 tails: tth tht htt So taking into account age exposes all the possibilities. Make sense?